OpenStudy (anonymous):

Plz Help!! Find the slope of a line on a graph (-3,0), (0,5)

OpenStudy (anonymous):

Slope of a line is rise over run so...\[\frac{ y-y_1 }{ x-x_1 }\]

OpenStudy (anonymous):

y=5/3x +5

OpenStudy (anonymous):

y=(5/3x) +5

OpenStudy (anonymous):

did you just need slope or the equation for the line?

OpenStudy (anonymous):

i need the slope

OpenStudy (anonymous):

5/3x is the slope

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

in the future, you can use the formula I gave you to figure it out. Knowing how to find the slope is more important than knowing the answer to a question. :)